A household bleach has a concentration of 4.60% (m/v) of sodium hypochlorite, NaOCl (aq). What mass of sodium hypochlorite does a 2.84 L container of this bleach contain.
1
Expert's answer
2012-12-12T08:23:27-0500
V = 2.84L - volume of container
w = 4.60% - percent of salt

mass is:
m = V*w/100%

m = 2.84*4.60% /100% = 0.13064 g